When a series of RLC circuit is subjected to 48V Vr=15V and Vl is 22V what is the voltage across the capacitor
1
Expert's answer
2020-11-06T10:12:38-0500

"V = \sqrt{V_r^2+(V_c – 22)^2}"

"48 = \sqrt{15^2+(V_c – 22)^2}"

"48^2 = 15^2 + (V_c \u2013 22)^2"

"V_c \u2013 22 = 45.59"

"V_c = 67.6 \;V"